i need help with the problem
-p/16 + 5 = 9

Answer by mickclns(59)

-p/16 + 5 = 9 Need to get -p/16 by itself, subtract 5 from each side (the opposite of adding 5)
-p/16 = 4 Need to get -p by itself, multiply both sides by 16 (the opposite of dividing by 16)
-p = 64 Need to multiply both sides by -1 (opposite of dividing by -1 or multiplying by -1)
p=-64 Solution